# Markarian 501 reincarnates with vigor as a temporary EHBL during VHE flaring in July 2014

### v1

@@ 3 VHE flaring of Markarian 501 during 2014
Mrk 501 was observed in multiwavelength by various instruments from July 16 to 31, 2014. During this period MAGIC telescopes observed for 13.5 h under dark and good atmospheric conditions (MAGIC Collaboration et al., 2020). The X-ray activity observed by the Neil Gehrels Swift Gamma-ray Burst Observatory was at its highest during its operational history and the spectra observed by different X-ray telescopes were very hard, similar to the large historic flare of 1997 when the synchrotron peak was shifted above 100 keV (Pian et al., 1998; Tavecchio et al., 2001), an indication of EHBL-like behavior. It is interesting to note that the flux variation during July 2014 in the energy regime covering the radio, optical and GeV energies was mild, but the enhanced emission in X-ray and VHE gamma-ray were significantly correlated ($>3\sigma$). Also, it was observed that the strength and the significance of the correlation increased for increasing energy in X-rays. For 0.3-2 keV the correlation was $\sim 3\sigma$ and for 2-10 keV the correlation was $\sim 4\sigma$ respectively. Temporal evaluation of the fluxes in X-ray and VHE gamma-ray were characterized on a day-by-day basis from july 16 to July 31, 2014 (MJD 56854–56869). Harder-when-brighter trends in X-ray spectra were observed, but the VHE spectra observed by MAGIC did not show such behavior. A summary of the VHE gamma-ray observations by MAGIC telescopes at different zenith angles is shown in Table 1 (MAGIC Collaboration et al., 2020).
@@
During the X-ray outburst, it was observed that the VHE gamma-ray flux above 0.15 TeV varied from $\sim 0.5$ to $\sim 2$ times the CU flux and no intra-night variability was observed in VHE gamma-rays during this two weeks flaring period. On 2014 July 19 (MJD56857.98) the highest X-ray flux above 2 keV was recorded by Swift while the MAGIC telescopes observed a rare unusual narrow spectral feature at $\sim 3$ TeV in the VHE spectrum. This intriguing peak-like feature is inconsistent with the standard analytic functions widely used to fit the VHE spectra of blazars. The observed broadband SED for 15 consecutive days, from July 16 to July 31, 2014, were explained using one-zone SSC models. Also, the narrow spectral feature on MJD 56857.98 is modeled by using the leptonic and the hadronic models (Hu & Yan, 2021; Petropoulou et al., 2024).
Table 1: The VHE observations of Mrk 501 by MAGIC telescopes during July 16 to July 31, 2014 are summarized in the table. The fourth column corresponds to the range of the observed zenith angle and the fifth column corresponds to the significance of the observation MAGIC Collaboration et al. (2020).
Date MJD Obs. time[h] Zd[] Significance
@@
Notes: A peak-like feature observed around 3 TeV.
Figure 1: The observed VHE spectra of Mrk 501 for the first three days are fitted using the two-zone photohadronic model. In all the figures, the normalization constants $F_{1}$ and $F_{2}$ are expressed in units of $10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$. The best fit to zone-1 is shown as the red curve and its extension as dotted red curve shows the behavior of the model at the high-energy limit if only zone-1 exists. Similarly, the spectrum in zone-2 is fitted with the black curve and the dashed black curve shows its behavior at the low-energy limit if only zone-2 exists. The dashed red and the dashed black curves are respectively the intrinsic curves corresponding to zone-1 and zone-2. The curves in Figures 2, 3, 4, and 5 have the same definitions and also, their normalization constants $F_{1}$ and $F_{2}$ are expressed in the same units as here.
Figure 2: The VHE spectrum of Mrk 501 observed on MJD 56857.98. See Figure 1 for the definitions of the curves.
Figure 3: The observed VHE spectra on MJD 56859.98, 56859.97, 56861.01 and 56862.02 are fitted with the two-zone photohadronic model. See Figure 1 for the definitions of the curves.
Figure 4: The VHE spectra observed on MJD 56863.00, 56864.02, 56865.00 and 56865.99 are fitted with the two-zone photohadronic model. See Figure 1 for the definitions of the curves.
Figure 5: The observed VHE spectra on MJD 56867.00, 56868.01, and 56869.93 are fitted with the two-zone photohadronic model. See Figure 1 for the definitions of the curves.
Figure 6: The cutoff energy $E^{c}_{\gamma}$ in the two-zone photohadronic model is shown for every flaring event. MJD 56857.98 corresponds to the day when a narrow peak-like feature at 3.18 TeV was observed and this is also shown in Figure 7, and in Figure 8.
Figure 7: The best fit values of the spectral indices $\delta_{1}$ and $\delta_{2}$ for every flaring event are shown.
4 Results
The Mrk501 is an extensively studied HBL in multiwavelength since its discovery and its VHE spectra are well explained by leptonic and hadronic models. However, during its VHE flaring in the years 1997, 2005 and 2012, it was observed that the synchrotron peak and the SSC peak were shifted towards higher energies. The synchrotron peaks were much above $10^{17}$ Hz, signaling that the emissions were temporary EHBL-like, thus presented difficulties to explain the VHE spectra using the leptonic model. Also, the standard photohadronic model was inadequate to explain these VHE SED. We use two-zone photohadronic model to explain these spectra well. The flaring of Mrk501 during July 2014 as shown in Table 1, had exactly the same behavior that of the flaring of 1997, 2005 and 2012, when the synchrotron peak was in the EHBL regime (above $10^{17}$ Hz) and the X-ray outburst was correlated with the VHE emission. So, we employ the two-zone photohadronic model to explain these 15 VHE spectra (from July 16 to July 31) on a day-by-day basis. A narrow peak-like feature at $\sim 3$ TeV was found for the first time on July 19, the day when the highest X-ray flux was also observed. This feature is inconsistent with the traditional interpretation of VHE spectra by using the different analytic functions. We shall also discuss in detail this narrow peak-like feature.
Figure 8: The best fit values of the normalization factors $F_{1}$ and $F_{2}$ in two-zone photohadronic model are shown for all the days.
@@ 4 Results
There were hour time difference between the VHE gamma-ray observations and the X-ray observations during the 15 consecutive days. Six observations have temporal difference of less than 1 h, five observations have time difference between 1 and 2 h, two observations have 3 h time difference and remaining two VHE observations do not have X-ray observations. However, during these 15 days observation period, there was no significant intra-night variability in the X-ray and the VHE gamma-ray bands. So, one can assume that the X-ray and the VHE gamma-ray observations were simultaneous.
@@
As stated above, the flaring of Mrk 501 during July 2014 was tEHBL-like, so, we shall use the two-zone photohadronic model to fit the VHE gamma-spectra. The VHE spectrum on MJD 56854.91 was observed in the energy range $0.13\,\mathrm{TeV}\leq E_{\gamma}\leq 1.3\,\mathrm{TeV}$ and is fitted well with $\delta_{1}=2.75$, $F_{1}=10.23\times 10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ for zone-1 and $\delta_{2}=3.70$, $F_{2}=2.70\times 10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ for zone-2 as shown in Figure 1. The cutoff energy between zone-1 and zone-2 is $E^{c}_{\gamma}=0.25$ TeV. For zone-1, $\delta_{1}=2.75$ corresponds to a high emission state and the intrinsic VHE photon flux $F_{\gamma,in}\propto E^{0.30}_{\gamma}$. But, for zone-2, $F_{\gamma,in}\propto E^{-0.70}_{\gamma}$ which is consistent with the fast fall of the flux above $E^{c}_{\gamma}$. Also, our fits to both the zones are very good. Similarly, on MJD 56855.91, the best fit to the spectrum is obtained for $\delta_{1}=2.80$, $F_{1}=11.50\times 10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ for zone-1 and $\delta_{2}=3.60$, $F_{2}=4.26\times 10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ for zone-2 as shown in Figure 1. Here the cutoff energy $E^{c}_{\gamma}=0.29$ TeV, slightly higher than the previous day. On the third day (MJD 56856.91), the zone-1 is in high emission state with $\delta_{1}=2.75$ and in zone-2, the flux decreases slower as compared to the previous day with $\delta_{2}=3.20$ and $E^{c}_{\gamma}=0.43$ TeV (Figure 1). It can be observed that $E^{c}_{\gamma}$ has increased slowly from 0.25 TeV, from the first day to 0.43 TeV, to the third day.
@@
A noticeable change in the spectrum on MJD 56857.98, on the fourth day was observed as shown in Figure 2. The spectrum was observed in the energy range $0.08\,\mathrm{TeV}\leq E_{\gamma}\leq 6.8\,\mathrm{TeV}$. Like previous days, zone-1 is still in the high emission state with $\delta_{1}=2.83$, but the cutoff energy drastically jumped to $E^{c}_{\gamma}=3.18$ TeV. Also, in zone-2, the spectrum falls extremely fast which can be fitted with $\delta_{2}=4.0$, which is shown in Figure 2. This is the first time we observe such a rapid fall of the VHE spectrum in zone-2 for tEHBL-like behavior of a HBL with a very large $\delta_{2}$ value. Previously, we had studied the tEHBL-like bahavior of Mrk 421, Mrk 501 and 1ES 1959+650 in different epochs using the two-zone photohadronic model, where we could explain the spectra very well and maximum value of $\delta_{2}$ was 3.5 (Sahu et al., 2017, 2018, 2020, 2021b). We had never observed such a fast fall of the spectrum in zone-2 and a cutoff energy very high. In this context, the VHE spectrum of Mrk 501 on MJD 56857.98 is unique and probably first of its type. Contrary to the claim by the MAGIC collaboration, our model does not predict a prominent peak at $E^{c}_{\gamma}=3.18$ TeV. We observe that the peak-like feature around $E^{c}_{\gamma}$ is due to the slight increase in the flux in zone-1 and then a rapid fall of the spectrum for $E_{\gamma}>E^{c}_{\gamma}$.
The VHE spectra from July 20 to July 24 (MJD 56858.98 to MJD 56862.02) are fitted very well which are shown in Figure 3. It can be seen that, the spectral index $\delta_{1}$ in zone-1 is always in the range $3.02.6<\delta_{1}<$ 3.0, corresponding to a high emission state. Also, in zone-2 we have $3.0<\delta_{2}\leq 3.5$ as observed in previous analysis to the VHE spectra of tEHBL-like behavior of several sources, including Mrk 501. On MJD 56858.98, the $E^{c}_{\gamma}=2.49$ TeV is high but the spectrum in zone-2 does not fall rapidly. So, we do not see any peak-like structure. On July 21 (MJD 56859.97), the cutoff energy drops down to very low, $E^{c}_{\gamma}=0.36$ TeV.
From July 25 to July 28 (MJD 56863.00 to MJD 56865.99), the spectra are fitted extremely well which are shown in Figure 4 with their respective parameters $F_{1}$, $F_{2}$, $\delta_{1}$ and $\delta_{2}$. Throughout the observation period, the maximum VHE photon energy observed is on MJD 56865.99, which is $E_{\gamma}=15.1$ TeV. In Figure 5, the VHE spectra of MJD 56867.00 to MJD 56869.93 (July 29 to July 31) are shown with their respective parameters. Here also, the spectra are fitted very well. On July 31, again the cutoff energy $E^{c}_{\gamma}=2.38$, which is high. However, the spectrum in zone-2 falls smoothly.
@@
We have shown the best fit values of $E^{c}_{\gamma}$ for all the 15 days observations in Figure 6. It can be seen that for the flaring on MJD 56857.98, the $E_{c}$ is the maximum and on next day it decreases slightly and then a free fall on MJD 56859.97 to $E^{c}_{\gamma}=0.36$ TeV. Also, we observe a oscillating behavior in the $E^{c}_{\gamma}$ values, which can be seen in Figure 6. Similarly, we have shown the best fit values of the VHE photon spectral indices $\delta_{1}$ and $\delta_{2}$ for zone-1 and zone-2 respectively in Figure 7. As the spectra in zone-1 are in high emission states, the values of $\delta_{1}$ are very similar for the whole period. However, in zone-2, we observe a variation in $\delta_{2}$ and again the largest value of $\delta_{2}=4.0$ is on MJD 56857.98.
@@
In Figure 8, the best fit values of the normalization factors $F_{1}$ and $F_{2}$ are plotted for the whole observation period. It can be seen that the variation in $F_{1}$ is minimal, however, there is a large variation in $F_{2}$ and again, on MJD 56857.98, the variation is the highest. These analysis clearly show that the VHE spectrum of MJD 56857.98 is very distinct from the rest.
@@
In the photohadronic model, high energy gamma-rays are produced from the interaction of high energy protons with the background SSC seed photons in the blazar jet. The region of the SSC flux responsible to produce the observed VHE gamma-rays lies in the low-energy tail region of the SSC spectrum which satisfies the condition in Eq. (1). For tEHBL-like behavior of an HBL, the VHE gamma-rays in zone-2 are produced when the accelerated high energy protons in the jet interact with the low energy MeV photons in the tail region of the SSC spectrum. The SSC flux in this region is $\Phi_{SSC}\propto\epsilon^{\beta_{2}}_{\gamma}\,\propto E^{-\beta_{2}}_{\gamma}$. Similarly, the SSC photon flux in zone-1 will behave as $\Phi_{SSC}\propto\epsilon^{\beta_{1}}_{\gamma}\,\propto E^{-\beta_{1}}_{\gamma}$. On MJD 56857.98, the best fit to zone-2 is achieved for $\delta_{2}=4.0$. With such a large $\delta_{2}$, the spectrum above $E^{c}_{\gamma}=3.18$ TeV falls rapidly. On this day, the highest energy photon observed is $E_{\gamma}=6.8$ TeV. This photon will be produced from the interaction of 68 TeV proton ($10\,E_{\gamma}$) with the SSC seed photon in the low energy tail region. Assuming that this seed photon has minimum energy $\epsilon_{\gamma}\simeq 7.5\times 10^{20}$ Hz ($\sim 3.1$ MeV)  (MAGIC Collaboration et al., 2020), the minimum bulk Lorentz factor is found to be $\Gamma\simeq 26.5$, which is consistent with the expectation. Using this $\Gamma$ we can estimate the $\epsilon^{c}_{\gamma}$ corresponding to the cutoff energy $E^{c}_{\gamma}=3.18$ TeV, which gives $\epsilon^{c}_{\gamma}\sim 6.6$ MeV. By adopting the proton spectral index $\alpha=2.0$ (Dermer & Schlickeiser, 1993), we get $\beta_{2}=2.0$ and $\beta_{1}=0.83$. The low energy tail region of the SSC flux (zone-2) behaves as $\Phi_{SSC}\propto\epsilon^{2}_{\gamma}$ in the energy range $3.1\,\mathrm{MeV}\lesssim\epsilon_{\gamma}\lesssim 6.6\,\mathrm{MeV}$ and for $\epsilon_{\gamma}\gtrsim 6.6\,\mathrm{MeV}$ but in zone-1, $\Phi_{SSC}\propto\epsilon^{0.83}_{\gamma}$. The increase of $\Phi_{SSC}$ is quadratic as a function of the seed photon energy $\epsilon_{\gamma}$ in zone-2 and implies that the intrinsic VHE flux falls as $F_{\gamma,in}\propto E^{-\alpha-\beta_{2}+3}_{\gamma}\propto E^{-1}_{\gamma}$. Similarly, in zone-1, the intrinsic flux increases mildly as $F_{\gamma,in}\propto E^{0.17}_{\gamma}$.

### v2

@@ 3 VHE flaring of Markarian 501 during 2014
Mrk 501 was observed in multiple wavelengths by various instruments from July 16 to 31, 2014. During this period, the MAGIC telescopes observed for 13.5 h under dark and good atmospheric conditions MAGIC Collaboration and others (2020b). The X-ray activity observed by the Neil Gehrels Swift Observatory was at its highest during its operational history and the spectra observed by different X-ray telescopes were very hard, similar to the large historic flare of 1997 when the synchrotron peak was shifted above 100 keV Pian and others (1998); Tavecchio and others (2001), an indication of tEHBL-like behavior. It is interesting to note that the flux variation during July 2014 in the energy regime covering the radio, optical and GeV energies was mild, but the enhanced emission in X-ray and VHE gamma-ray was significantly correlated ($3\sigma$). Also, it was observed that the strength and the significance of the correlation increased for increasing energy in X-rays. For 0.3-2 keV the correlation was $\sim 3\sigma$ and for 2-10 keV the correlation was $\sim 4\sigma$ respectively MAGIC Collaboration and others (2020b). Temporal evaluation of the fluxes in X-ray and VHE gamma-ray were characterized on a day-by-day basis from July 16 to July 31, 2014 (MJD 56854–56869). Harder-when-brighter trends in X-ray spectra were observed, but the VHE spectra observed by MAGIC did not show such behavior.
@@
During the X-ray outburst, it was observed that the VHE gamma-ray flux above 0.15 TeV varied from $\sim 0.5$ to $\sim 2$ times the CU flux and no intra-night variability was observed in VHE gamma-rays during this two weeks flaring period. On 2014 July 19 (MJD 56857.98) the highest X-ray flux above 2 keV was recorded by Swift, while the MAGIC telescopes observed a rare unusual narrow spectral feature at $\sim 3$ TeV in the VHE spectrum. This intriguing peak-like feature is inconsistent at more than $3\sigma$ with the standard analytic functions such as: Power Law, log-parabola and log-parabola with exponential cutoff MAGIC Collaboration and others (2020b), widely used to fit the VHE spectra of blazars. The observed broadband SED for 15 consecutive days, from July 16 to July 31, 2014, were explained using one-zone SSC models. Assuming that the narrow feature at $\sim 3$ TeV to be real, several theoretical scenarios were proposed to explain it MAGIC Collaboration and others (2020b). Also, this narrow spectral feature on MJD 56857.98 is modeled by using the leptonic and the hadronic models Hu and Yan (2021); Petropoulou et al. (2024).
Figure 1: The observed VHE spectra of Mrk 501 during the first three days are fitted with the two-zone photohadronic model. The normalization constants $F_{1}$ and $F_{2}$ are given in units of $10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$. The red and black solid curves represent the best fits for zone 1 and zone 2, respectively; dotted and dashed extension show their behaviors if each zone exist alone, The dashed red and black lines correspond to the intrinsic spectra of zone-1 and 2. The blue and salmon shaded regions indicate the $1\sigma$ confidence intervals for the best fits with ($F_{1}$, $\delta_{1}$) and ($F_{2}$, $\delta_{2}$), respectively. Figs. 2, 3, 4, and 5 follow the same definitions and units.
Figure 2: The VHE spectrum of Mrk 501 observed on MJD 56857.98. The red curve corresponds to zone-1 and the black curve to zone-2 (see Fig. 1 for definitions). Unlike Fig. 1, this observation shows a notably sharp flux suppression above $\sim$ 3 TeV, indicating an abrupt cutoff in the high-energy component. The gray dotted curve is the standard photohadronic model (one-zone) fit.
Figure 3: The observed VHE spectra on MJD 56858.98, 56859.97, 56861.01 and 56862.02 are fitted with the two-zone photohadronic model (see Fig. 1 for definitions). These spectra exhibit strong day-to-day variations in the cutoff energy. Additionally, none of the subfigures display a smooth or flat spectral shape, indicating significant dynamical changes in the background photon field during this period.
Figure 4: The VHE spectra observed on MJD 56863.00, 56864.02, 56865.00 and 56865.99 are fitted with the two-zone photohadronic model (see Fig. 1 for definitions). For MJD 56863.00, the last data point near $\sim$ 2.5 TeV should not be interpreted as a spectral peak, as it lacks a definite spectral index for zone-2.
Figure 5: The observed VHE spectra on MJD 56867.00, 56868.01, and 56869.93 are fitted with the two-zone photohadronic model (see Fig. 1 for definitions). Compared to previous days, the spectra in this period appear smoother, with a gradual increase in the cutoff energy $E^{c}_{\gamma}$. On MJD 56869.93, zone-2 has only two data points above $E^{c}_{\gamma}=2.38$ TeV. However, the spectrum in zone-2 is well described within the existing two-zone framework of the photohadronic model.
Figure 6: The cutoff energy $E^{c}_{\gamma}$ in the two-zone photohadronic model is shown for every flaring event. MJD 56857.98 corresponds to the day when a narrow peak-like feature at 3.18 TeV was observed and this is also shown in Fig. 7 (a) and Fig. 7 (b).
Figure 7: (a) The best fit values of the spectral indices $\delta_{1}$ and $\delta_{2}$ for every flaring event are shown. (b) The best fit values of the normalization factors $F_{1}$ and $F_{2}$ in two-zone photohadronic model are shown for all the days.
4 Results
The Mrk 501 is an extensively studied HBL in multiple wavelengths since its discovery and its VHE spectra are well explained by leptonic and hadronic models Abdo and others (2011); Furniss and others (2015); Ahnen and others (2017); MAGIC Collaboration and others (2020b); Mücke and Protheroe (2001); Mücke et al. (2003). However, during its VHE flaring in the years 1997, 2005 and 2012 Sahu et al. (2020), it was observed that the synchrotron peak and the SSC peak were shifted towards higher energies. The synchrotron peaks were much above $10^{17}$ Hz, signaling that the emissions were tEHBL-like, thus, presented difficulties in explaining the VHE spectra using the leptonic model. Also, the standard photohadronic model was inadequate to explain these VHE SED. We use two-zone photohadronic model to explain these spectra well. The flaring of Mrk 501 during July 2014 as shown in Table 1, had exactly the same behavior that of the flaring of 1997, 2005 and 2012, when the synchrotron peak was in the EHBL regime (above $10^{17}$ Hz) and the X-ray outburst was correlated with the VHE emission. So, we employ the two-zone photohadronic model to explain these 15 VHE spectra (from July 16 to July 31) on a day-by-day basis. A narrow peak-like feature at $\sim 3$ TeV was found for the first time on July 19, the day when the highest X-ray flux was also observed. This feature is inconsistent with the traditional interpretation of the VHE spectra using different analytic functions. We shall also discuss in detail this narrow peak-like feature.
Table 1: A summary of the parameters used for the calculation of the VHE spectra for all the days. The flux normalization factors $F_{1}$ and $F_{2}$ are in units of $10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$. The cutoff energy $E^{c}_{\gamma}$ is in TeV. The $1\sigma$ confidence level (CL) interval for the parameters $F_{1}$, $\delta_{1}$, $F_{2}$ and $\delta_{2}$ are shown.
MJD $F_{1}$ $\delta_{1}$ $F_{2}$ $\delta_{2}$ $E^{c}_{\gamma}$
@@ 4 Results
Notes: A peak-like feature observed around 3 TeV on MJD 56857.98.
@@
During these 15 consecutive days observation period, there were no strictly simultaneous observations in X-rays and gamma rays. However, there was no significant intra-night variability in the X-ray and the VHE gamma-ray bands. So, one can assume that the X-ray and the VHE gamma-ray observations were simultaneous.
@@
As stated above, the flaring of Mrk 501 during July 2014 was tEHBL-like, so, we shall use the two-zone photohadronic model to fit the VHE gamma-spectra. The VHE spectrum on MJD 56854.91 was observed in the energy range $0.13\,\mathrm{TeV}\leq E_{\gamma}\leq 1.3\,\mathrm{TeV}$ and is fitted well with $\delta_{1}=2.75$, $F_{1}=10.23\times 10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ for zone-1 and $\delta_{2}=3.70$, $F_{2}=2.70\times 10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ for zone-2 as shown in Fig. 1. The cutoff energy between zone-1 and zone-2 is $E^{c}_{\gamma}=0.25$ TeV. For zone-1, $\delta_{1}=2.75$ corresponds to a high emission state and the intrinsic VHE photon flux $F_{\gamma,in}\propto E^{0.30}_{\gamma}$. But, for zone-2, $\delta_{2}=3.70$ and $F_{\gamma,in}\propto E^{-0.70}_{\gamma}$ which is consistent with the fast fall of the flux above $E^{c}_{\gamma}$. Also, our fits to both zones are very good. Similarly, on MJD 56855.91, the best fit to the spectrum is obtained for $\delta_{1}=2.80$, $F_{1}=11.50\times 10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ for zone-1 and $\delta_{2}=3.60$, $F_{2}=4.26\times 10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ for zone-2 as shown in Fig. 1. Here the cutoff energy $E^{c}_{\gamma}=0.29$ TeV, slightly higher than the previous day. On the third day (MJD 56856.91), the zone-1 is in high emission state with $\delta_{1}=2.75$ and in zone-2, the flux decreases slower as compared to the previous day with $\delta_{2}=3.20$ and $E^{c}_{\gamma}=0.43$ TeV (Fig. 1). It can be observed that $E^{c}_{\gamma}$ has increased slowly from 0.25 TeV, from the first day to 0.43 TeV, to the third day. For each day, the value of $E^{c}_{\gamma}$ is adjusted depending on the behavior of the spectrum. The good fit values of the parameters $F_{1}$, and $\delta_{1}$ corresponding to zone-1 for each spectrum, the confidence interval at $1\sigma$ is shown in the blue shaded region. Similarly, for zone-2, the confidence interval at $1\sigma$ is shown in the salmon shaded region for good fit values of the parameters $F_{2}$, and $\delta_{2}$ for each spectrum.
A noticeable change in the spectrum was observed on MJD 56857.98, on the fourth day, as shown in Fig. 2. The VHE spectrum was observed in the energy range $0.08\,\mathrm{TeV}\leq E_{\gamma}\leq 6.8\,\mathrm{TeV}$. A peak-like feature is observed $\sim 3$ TeV and the spectrum falls rapidly above this energy, which is a new feature of this spectrum. Like previous days, we have used the two-zone photohadronic model to fit the spectrum. The zone-1 is still in the high emission state with $\delta_{1}=2.83$, but the cutoff energy has drastically jumped to $E^{c}_{\gamma}=3.18$ TeV. In the energy range $0.5\,\mathrm{TeV}\lesssim E_{\gamma}\lesssim 3.18\,\mathrm{TeV}$ (zone-1), the EBL corrected spectrum is attenuated with an inflection point at $E_{\gamma}\sim 1.82$ TeV.
In zone-2, the spectrum falls extremely fast, which can be fitted with $\delta_{2}=4.0$, as shown in Fig. 2. This is the first time we have observed such a rapid fall in the VHE spectrum in zone-2 for the tEHBL-like behavior of an HBL with a very large value of $\delta_{2}$. For $E^{c}_{\gamma}\,<\,3.18$ TeV the fitted VHE spectrum is in high emission state and increases slowly. However, for $E^{c}_{\gamma}\,>3.18$ TeV, the spectrum falls faster, resulting in a mild peak-like feature but not very prominent as claimed by the MAGIC Collaboration.
For comparison, we also fit the spectrum with the standard photohadronic model and the best fit values of the parameters are $\delta=3.0$ and $F_{0}=11.43\times 10^{-11}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$. This curve falls smoothly with no peak-like feature and in the high energy limit it falls slower than the observed spectrum (Fig. 2). Moreover, $\delta=3.0$ corresponds to low emission state and its intrinsic spectrum is flat which is common during a quiescent state of the blazar. This clearly shows that we can not use the standard photohadronic model (one-zone) to fit the spectrum of MJD 56857.98.
@@
Previously, we had studied the tEHBL-like behavior of Mrk 421, Mrk 501, and 1ES 1959+650 in different epochs using the two-zone photohadronic model, where we could explain the spectra very well and the maximum value of $\delta_{2}$ was 3.5 Sahu et al. (2020, 2021a, 2021b). We had never observed such a fast fall in the spectrum in zone-2 and a cutoff energy as high. In this context, the VHE spectrum of Mrk 501 on MJD 56857.98 is unique and probably the first of its type. The peak-like feature around $E^{c}_{\gamma}$ is observed to be due to a slight increase in the flux in zone-1 and then a rapid drop in the spectrum for $E_{\gamma}>E^{c}_{\gamma}$. The blue shaded region to the left and the salmon shaded region to the right in Fig. 2 correspond to the $1\sigma$ confidence interval for the good fit values of the parameters in zone-1 ($F_{1}$ and $\delta_{1}$), and in zone-2 ($F_{2}$ and $\delta_{2}$) respectively.
[…]
@@
We have shown the best fit values of $E^{c}_{\gamma}$ for all the 15 days observations in Fig. 6. It can be seen that for the flaring on MJD 56857.98, the $E^{c}_{\gamma}$ is the maximum and on next day it decreases slightly and then a free fall on MJD 56859.97 to $E^{c}_{\gamma}=0.36$ TeV. Also, we observe a oscillating behavior in the $E^{c}_{\gamma}$ values, which can be seen in Fig. 6. Similarly, we have shown the best fit values of the VHE photon spectral indices $\delta_{1}$ and $\delta_{2}$ for zone-1 and zone-2 respectively in Fig. 7(a). As the spectra in zone-1 are in high emission states, the values of $\delta_{1}$ are very similar for the whole period. However, in zone-2, we observe a variation in $\delta_{2}$ and again, the largest value of $\delta_{2}=4.0$ is on MJD 56857.98.
@@
In Fig. 7(b), the best fit values of the normalization factors $F_{1}$ and $F_{2}$ are plotted for the whole observation period. It can be seen that the variation in $F_{1}$ is minimal, however, there is a large variation in $F_{2}$ and again, on MJD 56857.98, the variation is the highest. These analyses clearly show that the VHE spectrum of MJD 56857.98 is very distinct from the rest. The best fit values of the parameters in the two-zone photohadronic model for all the 15 days observations are summarized in Table 1.
@@
On MJD 56857.98, the best fit to zone-2 is achieved for $\delta_{2}=4.0$. With such a large $\delta_{2}$, the spectrum above $E^{c}_{\gamma}=3.18$ TeV falls rapidly. On this day, the highest energy photon observed is $E_{\gamma}=6.8$ TeV. This photon will be produced from the interaction of 68 TeV proton ($10\,E_{\gamma}$) with the SSC seed photon in the low energy tail region. Assuming that this seed photon has minimum energy $\epsilon_{\gamma}\simeq 7.5\times 10^{20}$ Hz ($\sim 3.1$ MeV)  MAGIC Collaboration and others (2020b), the minimum bulk Lorentz factor can be estimated from Eq.(1) which is found to be $\Gamma\simeq 26.5$ and this is consistent with the expectation. Again, using this $\Gamma$ and the cutoff energy $E^{c}_{\gamma}=3.18$ TeV in Eq.(1), the corresponding cutoff energy for the SSC seed photon is $\epsilon^{c}_{\gamma}\sim 6.6$ MeV. By adopting the proton spectral index $\alpha=2.0$ Dermer and Schlickeiser (1993), we get $\beta_{2}=2.0$ and $\beta_{1}=0.83$. The low energy tail region of the SSC flux (zone-2) behaves as $\Phi_{SSC}\propto\epsilon^{2}_{\gamma}$ in the energy range $3.1\,\mathrm{MeV}\lesssim\epsilon_{\gamma}\lesssim 6.6\,\mathrm{MeV}$ and for $\epsilon_{\gamma}\gtrsim 6.6\,\mathrm{MeV}$ (zone-1), $\Phi_{SSC}\propto\epsilon^{0.83}_{\gamma}$. The increase of $\Phi_{SSC}$ is quadratic as a function of the seed photon energy $\epsilon_{\gamma}$ in zone-2 and implies that the intrinsic VHE flux falls as $F_{\gamma,in}\propto E^{-2.0-2.0+3}_{\gamma}\propto E^{-1}_{\gamma}$. Similarly, in zone-1, the intrinsic flux increases mildly as $F_{\gamma,in}\propto E^{-2.0-0.83+3}_{\gamma}\propto E^{0.17}_{\gamma}$.
The cutoff energy $E_{\gamma}^{c}$ divides the VHE spectrum into zone-1 and zone-2 with different photon spectral indices. Thus, for a given spectrum, the position of $E_{\gamma}^{c}$ in the two-zone photohadronic model is decided depending on the spectral behavior. If the spectrum shows a peak-like structure as in the case of MJD 56854.91 or MJD 56855.91 then the fitting is straightforward. However, in some cases the spectrum seems to be smooth and in these situations it is tricky to find the cutoff energy as in the case of MJD 56865.99. On MJD 56854.91 we do not have data above 1.3 TeV but a distinct peak at $\sim 0.25$ TeV and the two-zone photohadronic model fits well to the spectrum. Thus, the change in the spectral behavior is important to decide the position of the cutoff energy. Of course, availability of more data points in the high energy regime will help to improve the fit in zone-2, but not necessarily to fix the value of $E_{\gamma}^{c}$.